OpenStudy (anonymous):

HELLLLLLPPPPP number 1 is already done #1: Put the three functions in standard form. One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions. Create these three functions and explain to how you know these functions meet each condition. (Hint: Make sure that the b is even on g(x).) #2: Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. You may use the f(x) you created in question 1 as an example.

OpenStudy (anonymous):

#3: Find the solutions of g(x). Show each step. #4: Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences.

OpenStudy (anonymous):

1.f(x) = (x-3)(x-2) g(x) = (x - sqrt(2)(x-sqrt(3)) h(x) = (x-2i)(x-3i) i hope i did it right